Find the distance between the points (0,2) and (3,6)
Answer by sdmmadam@yahoo.com(530)

Find the distance between the points (0,2) and (3,6)
If A(x1,y1) and B(x2,y2)then by distance formula
we have (AB)^2 = [(x2-x1)^2 + (y2-y1)^2 ]
Given A(0,2), B(3,6)
(AB)^2 = [(x2-x1)^2 + (y2-y1)^2 ]
= [(3-0)^2 + (6-2)^2 ]
= [3^2 + 4^2]
=(9+16)
=25
(AB)^2 = 25
Therefore the distance AB between the two given points A(0,2) and B(3,6)is
AB = sqrt(25)= 5 units